A sample of h2o2 is x% by Mass x ml of kmno4 are required to oxidize one gram of this H2O2 sample calculate the molarity of kmno4 solution

Grade:11

1 Answers

Vikas TU
6 years ago
Dear Student,
The reaction is
              2 KmnO4+ 3 H2SO4 =  K2SO4 + 2 MnSO4 +3H2O+ 5[O]
              H2O2+ [O] à H2O + O2.   X5
              Hence 2 KmnO4 = 5 H2O2
              If the sample is of 1 gram KmnO4 reqired is =  2x158/ (5x34) =1.86 gram.
              So, normality is 1.86x10/31.6 = 0.58
              Molarity is 0.58/5 = 0.116.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)